# A Fourier-Jacobi Dirichlet series attached to modular forms of $\textup{SO}(2,4)$

### v1

@@ 10. Main Theorem
In this Section, we are finally ready to state and prove our main Theorem. The first step is to identify everything in the quaternion algebra $B$. We have already done that for the zeta functions $\zeta_{\xi_{i},\mathcal{P}}(s)$ in the previous Section. As we have already mentioned in Section 7, $\Xi(A^{+}(W))=B$. Moreover, $\Xi(A^{+}(\Lambda)\cap A^{+}(W))=\mathfrak{o}$ from [Shi06b, (2.19b)]. Therefore, $\Xi(A^{+}(\Lambda\cap W))=\mathfrak{o}$ from [Mur14, Lemma 2.1].
By substituting (30) to (15), we obtain
$\zeta_{\mathcal{P}}(2s-2k+4)D_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)\\ =L_{\mathcal{P}}(F;s-k+2)\sum_{i=1}^{h_{1}}\zeta_{\xi_{i},\mathcal{P}}(s-k+3)\sum_{j=1}^{h_{2}}A_{f_{j}}L_{\mathcal{P}}(\overline{f_{j}};s-k+5/2)^{-1}f_{j}(\tau(\zeta_{i}))=\\ =L_{\mathcal{P}}(F;s-k+2)\sum_{j=1}^{h_{2}}A_{f_{j}}L_{\mathcal{P}}(\overline{f_{j}};s-k+5/2)^{-1}\sum_{i=1}^{h_{1}}f_{j}(\tau(\zeta_{i}))\zeta_{\xi_{i},\mathcal{P}}(s-k+3).$ (35)
Note that as has been observed in [Psy25, Theorem 9.3], we have $L_{\textup{Sug}}\left(\textbf{1};s-1/2\right)=L_{\textup{Shi}}(s)$. Here, the notation means the $L$-function we encounter in [Sug85] and [Shi99] respectively, as these are normalised differently. From now on, we use the normalisation of Shimura. Therefore, from now on, we will write $L_{\mathcal{P}}(f_{j},s-k+3)$ instead of $L_{\mathcal{P}}(f_{j},s-k+5/2)$. However, we still keep the normalisation of Sugano for $L_{\mathcal{P}}(F,s)$.
@@
We set $S:=A^{+}(\Lambda\cap W)$ and $\mathcal{J}:=G^{+}(W)_{\mathbb{A}}\cap J$, and we select the finite set of primes $\mathcal{P}$ large enough so that $\mathcal{J}_{p}=(S_{p})^{\times}\cap G^{+}(W)_{p}$ for all $p\not\in\mathcal{P}$. We then consider the set $\mathcal{Y}:=\{y\in G^{+}(W)_{\mathbb{A}}\,\,|\,\,y_{p}\in\mathcal{Y}_{p}\textup{ for all }p\}$, where
$\mathcal{Y}_{p}=\begin{cases}G^{+}(W)_{p}\cap S_{p}\,\,\,\text{for}\,\,\,p\not\in\mathcal{P}\\ \mathcal{J}_{p}\,\,\text{otherwise}\end{cases}.$
We further define $\mathcal{Y}^{0}:=\{y\in\mathcal{Y}\,\,|\,\,\kappa(y)=\mathbb{Z}\}$, where $\kappa$ is defined in [Shi04, Section 22.8], and is called the capacity ideal. Moreover, we fix a set $\mathcal{A}\subset G^{+}(W)_{\mathbb{Q}}\backslash G^{+}(W)_{\mathbb{A}}/(G^{+}(W)_{\mathbb{A}}\cap J)$ of pre-images for the elements $\{v_{l}\}_{l=1}^{h_{2}}$ under $\tau$. Assume $\{a_{l}\}_{l=1}^{h_{2}}$ are these elements. We then define
@@
Moreover, we set $Q:=\left\{q\in\prod_{p}\mathbb{Z}_{p}\,\,|\,\,q_{p}\neq 0,\,\,q-1\in p\mathbb{Z}_{p},\,\forall p\in\mathcal{P}\right\}$, and $R:=\mathcal{J}\cap\mathbb{Q}^{\times}_{\mathbf{h}}$, where $\mathbb{Q}^{\times}_{\mathbf{h}}$ denotes the finite adeles of $\mathbb{Q}$. We then have the following Lemma.
@@
$\mathcal{Y}/\mathcal{J}=\bigsqcup_{\ell=1}^{h_{2}}\mathcal{Y}_{a_{\ell}}/\mathcal{J}$ (36)
@@
First of all, from [Shi04, Lemma 22.9, (iii)], we have that there is a bijection between the set $\mathcal{Y}/\mathcal{J}$ and the set $Q/R\times\mathcal{Y}^{0}/\mathcal{J}$ by mapping a $y\in\mathcal{Y}$ to $(q,y_{0})$, where $y=qy_{0}$ with $q\in Q$ and $y_{0}\in\mathcal{Y}^{0}$. Moreover, from [Shi04, Lemma 22.10] we have that
@@
Assume $K$ has class number $1$. We have the equality
$\sum_{\begin{subarray}{c}i\textup{ such that }\\ \rho(u_{i})=w_{\ell}\end{subarray}}\widehat{\zeta}_{\xi_{i},\mathcal{P}}(s)=\sum_{y\in\mathcal{Y}_{a_{\ell}}/\mathcal{J}}N(\nu(y)\mathbb{Z})^{-s},$
where $\nu$ is the map of (3) and $\rho$ the map of Proposition 7.2.
@@
In order to establish the equality, we first use the fact that the map $\Xi$ of Proposition 5.4 identifies $A^{+}(W)$ with $B$. Using this identification, we have that the map $\nu$ is the reduced norm of the quaternion algebra $B$ (see, for example, [Shi04, Section 7.3]). In particular, we may view the right-hand side of the claimed equality as a partial zeta series of $B$. Under this identification, we have that $\mathcal{Y}_{p}=\mathfrak{o}_{p}\cap B_{p}^{\times}$ for all $p\not\in\mathcal{P}$. Moreover, we have $\mathcal{J}_{p}=\mathfrak{o}_{p}^{\times}$ for all $p$ and hence in particular $\mathcal{Y}_{p}=\mathfrak{o}_{p}^{\times}$ for all $p\in\mathcal{P}$. Moreover the set $\mathcal{Y}_{a_{\ell}}/\mathcal{J}$ can be identified with locally principal right ideals in the order $\mathfrak{o}$ with the property that their image with respect to the map $\rho$ is the order corresponding to $w_{\ell}$ and locally are equal to $\mathfrak{o}_{p}$ for all $p\in\mathcal{P}$. Indeed, for an element $y=(y_{p})\in\mathcal{Y}$ we associate the locally principal ideal $I$ in $\mathfrak{o}$ such that $I_{p}=y_{p}\mathfrak{o}_{p}$ for all $p\not\in\mathcal{P}$, and $I_{p}=\mathfrak{o}_{p}$ for all $p\in\mathcal{P}$. By [Voi21, Main Theorem 16.7.7] such an ideal is invertible, and moreover every invertible (right) ideal in $\mathfrak{o}$ is locally principal. That is, the set $\mathcal{Y}_{a_{\ell}}/\mathcal{J}$ may be identified with the set of invertible right ideals in $\mathfrak{o}$ with the property that they are locally trivial for $p\in\mathcal{P}$, and their image with respect to the map $\rho$ is the order $w_{\ell}$. We now note that given an element $y=(y_{p})\in\mathcal{Y}$ corresponding to an ideal $I$, we have that $N(\nu(y)\mathbb{Z})=N(I)$. Indeed, by [Voi21, Main Theorem 16.7.7 (iv’)] and Lemma 9.4 above, we have that $N(I)=\textup{nrd}(I)$, where $\textup{nrd}(I)$ now equals $\textup{nrd}(I)=\textup{gcd}\{\nu(a)\,\,|\,\,a\in I\}$. Then, by [Voi21, Equation 16.3.4], we have that
@@
$\sum_{\mathcal{Y}_{a_{\ell}}/\mathcal{J}}N(\nu(y)\mathbb{Z})^{-s}=\sum_{\rho(\overline{\mathfrak{U}_{i}})=w_{\ell}}\sum_{\begin{subarray}{c}I\in[\overline{\mathfrak{U}_{i}}]\\ (I,{\mathcal{P}})=1\end{subarray}}N(I)^{-s},$
@@
We now define the space of automorphic forms on $G^{+}(W)_{\mathbb{A}}$ of weight $k$ with repsect to $\mathcal{J}$ as in [Shi04, (22.4a), (22.4b)]. We denote this space by $\mathcal{S}_{k}(\mathcal{J})$. Moreover, if $\psi$ is a Hecke character of $\mathbb{Q}$, we define
@@
as in [Shi04, (22.6)]. We also define the Hecke algebra $\mathfrak{R}(\mathcal{J},\mathcal{Y})$ in the sense of [Shi04, Section 8.16] and its action on $S_{k}(\mathcal{J,\psi})$ as follows. For an element $\tau\in\mathcal{Y}$, we take a finite subset $Y$ of $G^{+}(W)_{\mathbf{h}}$, so that
$\mathcal{J}\tau\mathcal{J}=\bigsqcup_{y\in Y}\mathcal{J}y.$ (37)
@@
$(\mathbf{g}\mid\mathcal{J}\tau\mathcal{J})(x):=\sum_{y\in Y}\mathbf{g}(xy^{-1}),$ (38)
@@
We now give the following Lemma, which is similar to [Shi04, (22.13)], but here we consider left instead of right coset representatives.
@@
First, as in [Shi04, p. 208], for each $\tau\in\mathcal{Y}$, we may find a set $Y_{\tau}$ such that
@@
Using now the definition of the action of (38), we have
@@
because $\mathbf{g}$ is of trivial character. Now we use the fact that $\mathcal{Y}^{*}=\mathcal{Y}$ since in our case $\mathcal{Y}_{p}=\mathfrak{o}_{p}$ for $p\not\in\mathcal{P}$, and $\mathfrak{o}_{p}^{\times}$ otherwise. But then we have that
@@
First we note that the eigenvalues of $\mathbf{g}$ are real. This follows from the fact that the Hecke operators are self-adjoint with respect to the corresponding inner product (see, for example, [Shi04, (22.11a)], with $\psi$ trivial there). Hence, if we write $\mu(\tau)\in\mathbb{R}$ for the Hecke eigenvalue of $g$ with respect to the Hecke operator $\mathcal{J}\tau\mathcal{J}$, with $\tau\in\mathcal{Y}$, and consider a decomposition as in (37), we have
@@
Assume $K$ has class number $1$ and it is not $\mathbb{Q}(\sqrt{-2})$. Choose $\ell$ so that it satisfies the conditions of Proposition 7.1. For each $j=1,\cdots,h_{2}$, we have
$\zeta_{K,\mathcal{P}}(s-k+3)\sum_{i=1}^{h_{1}}f_{j}(\tau(\zeta_{i}))\zeta_{\xi_{i},\mathcal{P}}(s-k+3)=f_{j}(1)L_{\mathcal{P}}\left(f_{j};s-k+3\right).$ (39)
@@
$\zeta_{K,\mathcal{P}}(s-k+3)\sum_{i=1}^{h_{1}}f_{j}(\tau(\zeta_{i}))\zeta_{\xi_{i},\mathcal{P}}(s-k+3)=\sum_{l=1}^{h_{2}}f_{j}(v_{l})\sum_{\begin{subarray}{c}i\textup{ such that }\\ \tau(\zeta_{i})=v_{l}\end{subarray}}\widehat{\zeta}_{\xi_{i},\mathcal{P}}(s-k+3).$
@@
$\sum_{\ell=1}^{h_{2}}f_{j}(\tau(a_{\ell}))\left(\sum_{\begin{subarray}{c}i\textup{ such that }\\ \rho(u_{i})=w_{\ell}\end{subarray}}\widehat{\zeta}_{\xi_{i},\mathcal{P}}(s-k+3)\right)=f_{j}(1)L_{\mathcal{P}}(g_{j};s-k+3).$ (40)
@@
Now, because $\tau$ is a homomorphism, we have $g_{j}(1)=f_{j}(\tau(1))=f_{j}(1)$. From (36) in Lemma 10.1, we can rewrite the above as
@@
It is now true from [Shi04, (22.12b)] that $\mathfrak{T}(s,g_{j})=L(g_{j},s)$. Proposition 10.2 now gives that
$\sum_{\begin{subarray}{c}i\textup{ such that }\\ \rho(u_{i})=w_{\ell}\end{subarray}}\widehat{\zeta}_{\xi_{i},\mathcal{P}}(s)=\sum_{y\in\mathcal{Y}_{a_{\ell}}/\mathcal{J}}N(\nu(y)\mathbb{Z})^{-s},$
and hence (40) follows. ∎
@@
Assume $K$ has class number $1$ and it is not $\mathbb{Q}(\sqrt{-2})$. Choose $\ell$ so that it satisfies the conditions of Proposition 7.1. Assume $F\in S_{k}(\Gamma_{S})$ is a Hecke eigenform, and consider the Poincaré series $\mathcal{P}_{k,D,r}$ of (10) with $D=-\ell q$ and $r\in L$ such that $(-\ell q,r)\in\widetilde{\textup{supp}}(L,\phi)$ of (9). Then, if $\mathcal{P}$ is a finite set of primes, including the primes described in [Psy25, Theorem 8.2], we have
$\zeta_{K,\mathcal{P}}(s-k+3)\zeta_{\mathcal{P}}(2s-2k+4)D_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)=A(\xi)L_{\mathcal{P}}(F;s-k+2),$
@@
By substituting (39) from Theorem 10.5 in (35), using also Lemma 10.4, we obtain
$\zeta_{K,\mathcal{P}}(s-k+3)\zeta_{\mathcal{P}}(2s-2k+4)D_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)=C\times L_{\mathcal{P}}(F;s-k+2),$
@@
We now write $\mathcal{V}_{k}$ for the span of $\mathcal{P}_{k,D,r}$ for all valid choices of $D$ and $r$, as in the Theorem above. We then obtain the following Corollary.
@@
Let $G\in\mathcal{V}_{k}$. Then, for a finite set of primes $\mathcal{P}$ sufficiently large we have that
$\zeta_{K,\mathcal{P}}(s-k+3)\zeta_{\mathcal{P}}(2s-2k+4)D_{F,G}(s)=\langle\phi_{1},\psi_{1}\rangle L_{\mathcal{P}}(F;s-k+2),$
where $\phi_{1}$ and $\psi_{1}$ are the first Fourier-Jacobi coefficients of $F$ and $G$, respectively.
@@
From the Theorem above, it follows that $\zeta_{K,\mathcal{P}}(s-k+3)\zeta_{\mathcal{P}}(2s-2k+4)D_{F,G}(s)$ is equal to $L_{\mathcal{P}}(F;s-k+2)$ up to a constant, when we take $\mathcal{P}$ large enough to apply our theorem for each Poincaré series in the linear combination that makes up $G$. This constant must then be $\langle\phi_{1},\psi_{1}\rangle$, by comparing the coefficients of $1^{-s}$ in the expansion of the Dirichlet series. ∎

### v2

@@ 12. Main Theorem
In this Section, we are finally ready to state and prove our main Theorem. The first step is to identify everything in the quaternion algebra $B$. We have already done that for the zeta functions $\zeta_{\mathcal{Q}}(\xi_{\mathfrak{m},i};s)$ in the previous Section. As we have already mentioned in Section 9, $\Xi(A^{+}(W))=B$. Moreover, $\Xi(A^{+}(\Lambda)\cap A^{+}(W))=\mathfrak{o}$ from [shimura2006quadratic, (2.19b)]. Therefore, $\Xi(A^{+}(\Lambda\cap W))=\mathfrak{o}$ from [murata2014, Lemma 2.1].
By substituting (8.7) to (11.2), we obtain
(12.1) $\zeta_{\mathcal{Q}}(2s-2k+4)D_{\mathcal{Q}}(\bm{F},\bm{P};s)\\ =L_{\mathcal{Q}}(F;s-k+2)\sum_{i=1}^{h_{1}}\widehat{\zeta}_{\mathcal{Q}}(\xi_{i};s-k+3)\sum_{j=1}^{h_{2}}A_{f_{j}}L_{\mathcal{Q}}(\overline{f_{j}};s-k+5/2)^{-1}f_{j}(\tau(\zeta_{i}))=\\ =L_{\mathcal{Q}}(F;s-k+2)\sum_{j=1}^{h_{2}}A_{f_{j}}L_{\mathcal{Q}}(\overline{f_{j}};s-k+5/2)^{-1}\sum_{i=1}^{h_{1}}f_{j}(\tau(\zeta_{i}))\widehat{\zeta}_{\mathcal{Q}}(\xi_{i};s-k+3).$
Note that as has been observed in [psyroukis_orthogonal, Theorem 9.3], we have $L_{\textup{Sug}}\left(\textbf{1};s-1/2\right)=L_{\textup{Shi}}(s)$. Here, the notation means the $L$-function we encounter in [sugano] and [shimura_orthogonal] respectively, as these are normalised differently. From now on, we use the normalisation of Shimura. Therefore, from now on, we will write $L_{\mathcal{Q}}(f_{j},s-k+3)$ instead of $L_{\mathcal{Q}}(f_{j},s-k+5/2)$. However, we still keep the normalisation of Sugano for $L_{\mathcal{Q}}(F,s)$.
@@
We set $S:=A^{+}(\Lambda\cap W)$ and $\mathcal{J}:=G^{+}(W)_{\mathbb{A}}\cap J$, and we select the finite set of primes $\mathcal{Q}$ large enough so that $\mathcal{J}_{p}=(S_{p})^{\times}\cap G^{+}(W)_{p}$ for all $p\not\in\mathcal{Q}$. We then consider the set $\mathcal{Y}:=\{y\in G^{+}(W)_{\mathbb{A}}\,\,|\,\,y_{p}\in\mathcal{Y}_{p}\textup{ for all }p\}$, where
$\mathcal{Y}_{p}=\begin{cases}G^{+}(W)_{p}\cap S_{p}\,\,\,\text{for}\,\,\,p\not\in\mathcal{Q}\\ \mathcal{J}_{p}\,\,\text{otherwise}\end{cases}.$
We further define $\mathcal{Y}^{0}:=\{y\in\mathcal{Y}\,\,|\,\,\kappa(y)=\mathbb{Z}\}$, where $\kappa$ is defined in [shimura2004arithmetic, Section 22.8], and is called the capacity ideal. Moreover, we fix a set $\mathcal{A}\subset G^{+}(W)_{\mathbb{Q}}\backslash G^{+}(W)_{\mathbb{A}}/(G^{+}(W)_{\mathbb{A}}\cap J)$ of pre-images for the elements $\{v_{l}\}_{l=1}^{h_{2}}$ under $\tau$. Assume $\{a_{l}\}_{l=1}^{h_{2}}$ are these elements.
We then define
@@
Moreover, we set $Q:=\left\{q\in\prod_{p}\mathbb{Z}_{p}\,\,|\,\,q_{p}\neq 0,\,\,q-1\in p\mathbb{Z}_{p},\,\forall p\in\mathcal{Q}\right\}$, and $R:=\mathcal{J}\cap\mathbb{Q}^{\times}_{\mathbf{h}}$, where $\mathbb{Q}^{\times}_{\mathbf{h}}$ denotes the finite adeles of $\mathbb{Q}$. We then have the following Lemma.
@@
(12.2) $\mathcal{Y}/\mathcal{J}=\bigsqcup_{\ell=1}^{h_{2}}\mathcal{Y}_{a_{\ell}}/\mathcal{J}$
@@
First of all, from [shimura2004arithmetic, Lemma 22.9, (iii)], we have that there is a bijection between the set $\mathcal{Y}/\mathcal{J}$ and the set $Q/R\times\mathcal{Y}^{0}/\mathcal{J}$ by mapping a $y\in\mathcal{Y}$ to $(q,y_{0})$, where $y=qy_{0}$ with $q\in Q$ and $y_{0}\in\mathcal{Y}^{0}$. Moreover, from [shimura2004arithmetic, Lemma 22.10] we have that
@@
We have the equality
$\sum_{\begin{subarray}{c}i\textup{ such that }\\ \rho(u_{i})=w_{\ell}\end{subarray}}\widehat{\zeta}_{\mathcal{Q}}(\xi_{i};s)=\sum_{y\in\mathcal{Y}_{a_{\ell}}/\mathcal{J}}N(\nu(y)\mathbb{Z})^{-s},$
where $\nu$ is the map of (2.2) and $\rho$ the map of Proposition 9.3.
@@
In order to establish the equality, we first use the fact that the map $\Xi$ of Proposition 7.4 identifies $A^{+}(W)$ with $B$. Using this identification, we have that the map $\nu$ is the reduced norm of the quaternion algebra $B$ (see, for example, [shimura2004arithmetic, Section 7.3]). In particular, we may view the right-hand side of the claimed equality as a partial zeta series of $B$. Under this identification, we have that $\mathcal{Y}_{p}=\mathfrak{o}_{p}\cap B_{p}^{\times}$ for all $p\not\in\mathcal{Q}$. Moreover, we have $\mathcal{J}_{p}=\mathfrak{o}_{p}^{\times}$ for all $p$ and hence in particular $\mathcal{Y}_{p}=\mathfrak{o}_{p}^{\times}$ for all $p\in\mathcal{Q}$. Moreover the set $\mathcal{Y}_{a_{\ell}}/\mathcal{J}$ can be identified with locally principal right ideals in the order $\mathfrak{o}$ with the property that their image with respect to the map $\rho$ is the order corresponding to $w_{\ell}$ and locally are equal to $\mathfrak{o}_{p}$ for all $p\in\mathcal{Q}$. Indeed, for an element $y=(y_{p})\in\mathcal{Y}$ we associate the locally principal ideal $I$ in $\mathfrak{o}$ such that $I_{p}=y_{p}\mathfrak{o}_{p}$ for all $p\not\in\mathcal{Q}$, and $I_{p}=\mathfrak{o}_{p}$ for all $p\in\mathcal{Q}$. By [voight, Main Theorem 16.7.7] such an ideal is invertible, and moreover every invertible (right) ideal in $\mathfrak{o}$ is locally principal. That is, the set $\mathcal{Y}_{a_{\ell}}/\mathcal{J}$ may be identified with the set of invertible right ideals in $\mathfrak{o}$ with the property that they are locally trivial for $p\in\mathcal{Q}$, and their image with respect to the map $\rho$ is the order $w_{\ell}$. We now note that given an element $y=(y_{p})\in\mathcal{Y}$ corresponding to an ideal $I$, we have that $N(\nu(y)\mathbb{Z})=N(I)$. Indeed, by [voight, Main Theorem 16.7.7 (iv’)] and Lemma 11.4 above, we have that $N(I)=\textup{nrd}(I)$, where $\textup{nrd}(I)$ now equals $\textup{nrd}(I)=\textup{gcd}\{\nu(a)\,\,|\,\,a\in I\}$. Then, by [voight, Equation 16.3.4], we have that
@@
$\sum_{\mathcal{Y}_{a_{\ell}}/\mathcal{J}}N(\nu(y)\mathbb{Z})^{-s}=\sum_{\rho(\overline{\mathfrak{U}_{i}})=w_{\ell}}\sum_{\begin{subarray}{c}I\in[\overline{\mathfrak{U}_{i}}]\\ (I,{\mathcal{Q}})=1\end{subarray}}N(I)^{-s},$
@@
We now define the space of automorphic forms on $G^{+}(W)_{\mathbb{A}}$ of weight $k$ with repsect to $\mathcal{J}$ as in [shimura2004arithmetic, (22.4a), (22.4b)]. We denote this space by $\mathcal{S}_{k}(\mathcal{J})$. Moreover, if $\psi$ is a Hecke character of $\mathbb{Q}$, we define
@@
as in [shimura2004arithmetic, (22.6)]. We also define the Hecke algebra $\mathfrak{R}(\mathcal{J},\mathcal{Y})$ in the sense of [shimura2004arithmetic, Section 8.16] and its action on $S_{k}(\mathcal{J,\psi})$ as follows. For an element $\tau\in\mathcal{Y}$, we take a finite subset $Y$ of $G^{+}(W)_{\mathbf{h}}$, so that
(12.3) $\mathcal{J}\tau\mathcal{J}=\bigsqcup_{y\in Y}\mathcal{J}y.$
@@
(12.4) $(\mathbf{g}\mid\mathcal{J}\tau\mathcal{J})(x):=\sum_{y\in Y}\mathbf{g}(xy^{-1}),$
@@
We now give the following Lemma, which is similar to [shimura2004arithmetic, (22.13)], but here we consider left instead of right coset representatives.
@@
First, as in [shimura2004arithmetic, p. 208], for each $\tau\in\mathcal{Y}$, we may find a set $Y_{\tau}$ such that
@@
Using now the definition of the action of (12.4), we have
@@
because $\mathbf{g}$ is of trivial character.
Now we use the fact that $\mathcal{Y}^{*}=\mathcal{Y}$ since in our case $\mathcal{Y}_{p}=\mathfrak{o}_{p}$ for $p\not\in\mathcal{Q}$, and $\mathfrak{o}_{p}^{\times}$ otherwise. But then we have that
@@
First we note that the eigenvalues of $\mathbf{g}$ are real. This follows from the fact that the Hecke operators are self-adjoint with respect to the corresponding inner product (see, for example, [shimura2004arithmetic, (22.11a)], with $\psi$ trivial there). Hence, if we write $\mu(\tau)\in\mathbb{R}$ for the Hecke eigenvalue of $g$ with respect to the Hecke operator $\mathcal{J}\tau\mathcal{J}$, with $\tau\in\mathcal{Y}$, and consider a decomposition as in (12.3), we have
@@
Assume $K=\mathbb{Q}(\sqrt{-m})$ with $m$ square free and $m\not\equiv 2\pmod{4}$. Choose $\ell$ so that it satisfies the conditions of Proposition 9.1. For each $j=1,\cdots,h_{2}$, we have
(12.5) $\sum_{i=1}^{h_{1}}f_{j}(\tau(\zeta_{i}))\widehat{\zeta}_{\mathcal{Q}}(\xi_{i};s-k+3)=f_{j}(1)L_{\mathcal{Q}}\left(f_{j};s-k+3\right).$
@@
$\sum_{i=1}^{h_{1}}f_{j}(\tau(\zeta_{i}))\widehat{\zeta}_{\mathcal{Q}}(\xi_{i};s-k+3)=\sum_{l=1}^{h_{2}}f_{j}(v_{l})\sum_{\begin{subarray}{c}i\textup{ such that }\\ \tau(\zeta_{i})=v_{l}\end{subarray}}\widehat{\zeta}_{\mathcal{Q}}(\xi_{i};s-k+3).$
@@
(12.6) $\sum_{\ell=1}^{h_{2}}f_{j}(\tau(a_{\ell}))\left(\sum_{\begin{subarray}{c}i\textup{ such that }\\ \rho(u_{i})=w_{\ell}\end{subarray}}\widehat{\zeta}_{\mathcal{Q}}(\xi_{i};s-k+3)\right)=f_{j}(1)L_{\mathcal{Q}}(g_{j};s-k+3).$
@@
Now, because $\tau$ is a homomorphism, we have $g_{j}(1)=f_{j}(\tau(1))=f_{j}(1)$. From (12.2) in Lemma 12.1, we can rewrite the above as
@@
It is now true from [shimura2004arithmetic, (22.12b)] that $\mathfrak{T}(s,g_{j})=L(g_{j},s)$. Proposition 12.2 now gives that
$\sum_{\begin{subarray}{c}i\textup{ such that }\\ \rho(u_{i})=w_{\ell}\end{subarray}}\widehat{\zeta}_{\mathcal{Q}}(\xi_{i};s)=\sum_{y\in\mathcal{Y}_{a_{\ell}}/\mathcal{J}}N(\nu(y)\mathbb{Z})^{-s},$
and hence (12.6) follows. ∎
@@
Assume $K=\mathbb{Q}(\sqrt{-m})$ with $m$ square free and $m\not\equiv 2\pmod{4}$. Choose $\ell$ so that it satisfies the conditions of Proposition 9.1. Then we have,
$\zeta_{\mathcal{Q}}(2s-2k+4)D_{\mathcal{Q}}(\bm{F},\bm{P};s)=A(\xi)L_{\mathcal{Q}}(F;s-k+2).$
@@
By substituting (12.5) from Theorem 12.5 in (12.1), using also Lemma 12.4, we obtain
$\zeta_{\mathcal{Q}}(2s-2k+4)D_{\mathcal{Q}}(\bm{F},\bm{P};s)=C\times L_{\mathcal{Q}}(F;s-k+2),$
@@
[…]
@@
Let $\bm{G}\in\bm{V}_{k}$. Then, for a finite set of primes $\mathcal{Q}$ sufficiently large we have that
$\zeta_{\mathcal{Q}}(2s-2k+4)D_{\mathcal{Q}}(\bm{F},\bm{G};s)=\widetilde{C}\,L_{\mathcal{Q}}(F;s-k+2),$
where $\widetilde{C}\in\mathbb{C}$ is given as
$\widetilde{C}=(\#\textup{Cl}(K/\mathbb{Q}))\sum_{\mathfrak{m}\in\textup{Cl}(K)/\textup{Cl}(K/\mathbb{Q})}\langle\phi_{\mathfrak{m},1},\psi_{\mathfrak{m},1}\rangle.$
Here $\phi_{\mathfrak{m},1}$ and $\psi_{\mathfrak{m},1}$ are the first Fourier-Jacobi coefficients of $F_{\mathfrak{m}}$ and $G_{\mathfrak{m}}$, respectively.
@@
From the Theorem above, it follows that $\zeta_{\mathcal{Q}}(2s-2k+4)D_{\mathcal{Q}}(\bm{F},\bm{G};s)$ is equal to $L_{\mathcal{Q}}(F;s-k+2)$ up to a constant, when we take $\mathcal{Q}$ large enough to apply our Theorem for each Poincaré series in the linear combination that makes up $G$. This constant is equal to $\widetilde{C}$, by comparing the coefficients of $1^{-s}$ in the expansion of the Dirichlet series. ∎